OpenStudy (anonymous):

For what x-value(s) does sin(x) = -1? A. -270°, -90°, 90°, and 270° B. -270°, -90°, 0°, 90°, and 270° C. -90 and 270° D. -270° and 90°

OpenStudy (anonymous):

C. -90 and 270

OpenStudy (anonymous):

sin= y/r -90= (0, -1) 270= ( 0,-1) y/r= -1/1=-1 y/r= -1/1 = -1